OpenStudy (anonymous):

can anyone talk with me about improper integrals? i have a question or two and it should only be a few mins

OpenStudy (anonymous):

When you are given an integral how do know when to test if it converges or diverges?'

OpenStudy (amistre64):

when to test? usually when you are asked: does this converge or diverge :/

OpenStudy (anonymous):

ok so i have some questons where its the integral from 0 to infinity. i know to add the limit part when there is an infinity there. But when it is not that obvious when do iadd it? i think vertical asymptotes?

OpenStudy (amistre64):

an integral is more compact than discrete stuff; so if the integral version goes to zero; the discrete does too

OpenStudy (amistre64):

if i had a problem to work on it be easier to demonstarte

OpenStudy (amistre64):

me no type gooder this nite lol

OpenStudy (anonymous):

ok so for example integral from 0 to infinity of dx/(SQRT(x)(x+4))

OpenStudy (anonymous):

so i took that and because there is an infinity a add the lim as b approces inf of the integral from 0 to b

OpenStudy (amistre64):

\[\int_{0}^{inf}\ \frac{1}{(x+4)\sqrt{x}}dx\]

OpenStudy (amistre64):

this perchance?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

thats bad at zero as well

OpenStudy (anonymous):

see thats what i thought but in the key i got from the professor he only added the one that approaches infinity

OpenStudy (amistre64):

the infinity part just makes the bottom huge and this thing goes to 0; but for propercalities ...

OpenStudy (amistre64):

-4 and 0 are bad

OpenStudy (amistre64):

so we need to split this up into prolly 4 integrals to test out

OpenStudy (anonymous):

ahhh ok good so im not crazy lol

OpenStudy (anonymous):

you dont need to do the work for it its alright i just wanted to make sure i was not under the wrong impression

OpenStudy (amistre64):

\[\int_{t}^{1}\frac{1}{(x+4)\sqrt{x}}dx+\int_{1}^{u}\frac{1}{(x+4)\sqrt{x}}dx+\int_{u}^{10}\frac{1}{(x+4)\sqrt{x}}dx +\int_{10}^{w}\frac{1}{(x+4)\sqrt{x}}dx\]

OpenStudy (amistre64):

i figure 1 and 10 were good numbers; just need something to measure form so the values themselves just need to be inside the good interval parts

OpenStudy (amistre64):

lol... helps if i read; 0 to inf so -4 aint even in the scope

OpenStudy (anonymous):

ok i see that makes sense. So when im solving any integral problem is the first thing i should check is if it has vertical asymptotes? because if it does i need to do the limit parts right

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

right; and in this case -4 is a VA but its not even in our interval to begin with so 0 and inf are the trouble spots

OpenStudy (anonymous):

ok makes sense so im just trying to figure out the thought process for integrals

OpenStudy (amistre64):

find a good spot to measure from; then limit it out to the trouble spots

OpenStudy (anonymous):

what it was always before was Integration by parts, Partial deriv, trig sub or u-sub then solve. but now that we have learned this the first step should be to find vertical asymptotes and split it up then figure out what method to use for integration?

OpenStudy (amistre64):

yep, sounds planish to me

OpenStudy (anonymous):

ok cool. actually, looking back does it matter? if you find the antideriv and split up by vertical asymptotes before you do the interval you would get the same answer it looks like

OpenStudy (amistre64):

the asymps just help to define your plan of attack; so get comfortable with it in any manner you want

OpenStudy (amistre64):

since the same curve is being analysed throught out; then yeah, you can integrate first and get it set up for the limit processes

OpenStudy (anonymous):

wow i guess i just needed to write it all out to realize it lol. TY forhelping and...listening? lol and making sure i dident say anyting dumb

OpenStudy (amistre64):

yw :) and good luck

OpenStudy (anonymous):

u too